What is the mass of ethanol that must be combusted to generate 20.4 L of CO2 at STP?


1
Expert's answer
2022-03-31T09:50:04-0400

C2H5OH + 3O2 → 2CO2 + 3H2O

20.4L = 20.4kg= 20400g

Moles of CO2 produced= 20400/44= 463.6

Moles of ethanol that burn= 463.6/2= 231.8

Mass of ethanol= 231.8× 46= 10663.64g